OpenStudy (anonymous):

How do you solve the algebraic equation (16/81)^-1/3

OpenStudy (chaise):

(16/81)^-1/3 = 1/(16/81)^1/3 = \[1/\sqrt[3]{16/81}^1\] Need any further explanation?

OpenStudy (anonymous):

you can easily solve by logirthm

OpenStudy (anonymous):

The answer isn't the problem I know the answer just not how to get the answer

OpenStudy (anonymous):

in fact it is not an equation, it is a number. the number is \[(\frac{81}{16})^{\frac{1}{3}}\]

OpenStudy (anonymous):

rather it means \[\frac{\sqrt[3]{81}}{\sqrt[3]{16}}\]

OpenStudy (anonymous):

since \[16=2^4\] and \[81=3^4\] you can rewrite this as \[\frac{3\sqrt{3}}{2\sqrt{2}}\]

OpenStudy (anonymous):

dang i am making mistake after mistake. i meant it is \[\frac{3\sqrt[3]{3}}{2\sqrt[3]{2}}\]

OpenStudy (anonymous):

Some how it is supposed to be 3.375

OpenStudy (anonymous):

then it is a calculator exercise

OpenStudy (anonymous):

at satellite73 the problem actually is how to find 2^(1/3)

OpenStudy (anonymous):

take the cubed root of \[\frac{81}{16}\] and see what you get

OpenStudy (anonymous):

Not supposed to be

OpenStudy (anonymous):

and the method is similar as to find out square root of 3

OpenStudy (anonymous):

newton's method? is that what you are using?

OpenStudy (anonymous):

which newton's method are you thinking

OpenStudy (anonymous):

the method I'm talking it's the method that I learnt in primary class.

OpenStudy (anonymous):

and I don't know who gave this